OpenStudy (anonymous):

Find the current that flows through the filament of a 400 watts flat iron connected to a 220 Volt power line ( 50 mA, 5 A, 5mA, 500 mA )

OpenStudy (anonymous):

Use: \[P=\frac{ V^2 }{ R }\] To find the resistance and then use \[V = IR\] to find the current

OpenStudy (anonymous):

already try that formula but it didnt help to find the correct answer... oh yeah forgot to insert the choices... ( 50 mA, 5 A, 5mA, 500 mA )

OpenStudy (gorv):

have type the Q completely??

OpenStudy (anonymous):

thats all the question... what really bothers me if there's a constant that you have to use in it...

OpenStudy (anonymous):

The answer I am getting is : \(I = 1.81 \; Amp\)

OpenStudy (anonymous):

Using net, I have downloaded this Document, check \(Ques : 34\) of the Document :

OpenStudy (gorv):

we use P=VI 400=220*i i=400/220=1.8 amp

OpenStudy (gorv):

so this is the max current can flow through the device

OpenStudy (gorv):

500ma=0.5 amp is most suited ans to this

OpenStudy (gorv):

@entrial its correct u can try buddy i m sure :)

OpenStudy (unklerhaukus):

some of the voltage drop will be in the iron but, not at the filament

OpenStudy (anonymous):

@gorv yes 0.5 A or 500mA is the correct answer just want to know how did he get that answer and @waterineyes were also getting the same output...

OpenStudy (gorv):

yeah bcozz it is the nearest value

OpenStudy (gorv):

we can consider that some current will flow in other part

OpenStudy (gorv):

more than 1.8 A will not flow so nearest to 1.8 is 0.5

OpenStudy (anonymous):

ok @gorv your the BAT thanks....

OpenStudy (gorv):

welcme buddy :)